OpenStudy (anonymous):

Math question: Write the equation of the line with an undefined slope, passing through the point (2, 5).

OpenStudy (anonymous):

Equation with undefined slope and pint (x1, y1) is: \[y - y _{1} = m(x - x _{1})\] Where "m" is that undefined slope.

OpenStudy (anonymous):

So do I write the equation like y - 5 = m(x-2)

OpenStudy (anonymous):

Yes, you have got it. You can write that equation in a simpler form, though.

OpenStudy (anonymous):

y = m(x-2) + 5 gives y as a function of x and is almost slope-intercept form.

OpenStudy (anonymous):

y = mx + (5 - 2m) is actually slope-intercept if we consider (5 - 2m) as the y-intercept (which it actually is) though this intercept is in terms of "m".

OpenStudy (anonymous):

Okay, thank you